OpenStudy (anonymous):

1) When the sum of 5 and twice a positive number is subtracted from the square of the number, 0 results. Find the number.

pooja195 (pooja195):

Let x be the number First thing it says is subtracted from the square of A NUMER (x) sooo \[\huge~\rm~x^2- \]

pooja195 (pooja195):

*number

pooja195 (pooja195):

Then it says the number is twice postivie soo 2x and the sum (adding) of 5 so it would look like this \[\huge~\rm~(2x+5)\]

pooja195 (pooja195):

The result is 0 so =0 \[\huge~\rm~x^2-(2x+5)=0 \]

pooja195 (pooja195):

We can remove the parentheses and write it like this: \[\huge~\rm~x^2-2x-5=0 \]

pooja195 (pooja195):

Find the abc values :) a ,b , c values are \[\huge\rm Ax^2+Bx+C=0\] where a =leading coefficient b= leading coefficient c= constant term

OpenStudy (anonymous):

a=1 b=2 c=5

OpenStudy (anonymous):

-2 -5

pooja195 (pooja195):

Good : )

pooja195 (pooja195):

\[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[x=\frac{ 2\pm \sqrt{2^2-4(1)(5)} }{ 2(1) }\]

pooja195 (pooja195):

\[\huge~\rm~x=\frac{ -(-2) \pm \sqrt{(-2)^2-4(1)(-5)} }{ 2(1) }\]

OpenStudy (anonymous):

oh in the sqrt its -2

OpenStudy (anonymous):

and -5 okay

pooja195 (pooja195):

ok can you solve whats in the sqrt?

OpenStudy (anonymous):

\[\sqrt{80}\]

OpenStudy (anonymous):

\[2\pm4\sqrt{5}\]

pooja195 (pooja195):

I got \[\huge~\rm~1±1\sqrt{6}\]

OpenStudy (anonymous):

..how?

pooja195 (pooja195):

eh i cheated and used a calc ;p heres the work http://prntscr.com/80bm1y

OpenStudy (anonymous):

what calc did you use

pooja195 (pooja195):

Then we make 2 seperate equations one plus and one minus since we basically have to split the plus minus sign then u simplfy from here \[\frac{ \huge~\rm~1+1\sqrt{6} }{ \huge~\rm~1 }\] \[\frac{ \huge~\rm~1-1\sqrt{6} }{ \huge~\rm~1 }\]

pooja195 (pooja195):

You will end up with ONE positive answer and ONE negative answer. The psotive answer is the answer to the question :)

OpenStudy (anonymous):

...

pooja195 (pooja195):

...?

OpenStudy (anonymous):

how

pooja195 (pooja195):

I never learned....:/ you will need to use a calculator or ask someone else...i have always used a calc for this part... @Astrophysics

OpenStudy (ali2x2):

@ganeshie8

OpenStudy (just_one_last_goodbye):

@pooja195 what do we want to convert into a decimal?

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Then we make 2 seperate equations one plus and one minus since we basically have to split the plus minus sign then u simplfy from here \[\frac{ \huge~\rm~1+1\sqrt{6} }{ \huge~\rm~1 }\] \[\frac{ \huge~\rm~1-1\sqrt{6} }{ \huge~\rm~1 }\] \(\color{#0cbb34}{\text{End of Quote}}\) This and show your stpes

pooja195 (pooja195):

*steps

OpenStudy (just_one_last_goodbye):

alright

OpenStudy (just_one_last_goodbye):

both on top of each other like fractions correct?

OpenStudy (ali2x2):

@just_one_last_goodbye i think so

pooja195 (pooja195):

Actually a better way to solve it Apply this rule \[\huge~\rm~\frac{ a }{ 1 }=a\]

pooja195 (pooja195):

\[\huge~\rm~\frac{1+\sqrt{6}1}{1}=1+\sqrt{6}1 \]

pooja195 (pooja195):

\[\huge~\rm~1+1\cdot \sqrt{6} \]

pooja195 (pooja195):

\[\huge~\rm~1\cdot \:a=a\]

pooja195 (pooja195):

\[\huge~\rm~1+\sqrt{6}\]

pooja195 (pooja195):

Now turn it into decimal form

pooja195 (pooja195):

We dont even need to solve fore the subtraction problem because we know that the number will be negative

OpenStudy (anonymous):

ok

OpenStudy (ali2x2):

ill try it myself

OpenStudy (just_one_last_goodbye):

I got 3.44949

pooja195 (pooja195):

Thats right

OpenStudy (anonymous):

I was gonna ask if we round it to 3.45

pooja195 (pooja195):

Yep :)

OpenStudy (ali2x2):

i got 3.44949

OpenStudy (anonymous):

okay, thank you for being so patient.

OpenStudy (ali2x2):

oh they already said that

pooja195 (pooja195):

lol

OpenStudy (ali2x2):

i cri